# Fine-Tuned Machine Translation Metrics Struggle in Unseen Domains

### v1

@@ Abstract
We introduce a new, extensive multidimensional quality metrics (MQM) annotated dataset covering 11 language pairs in the biomedical domain. We use this dataset to investigate whether machine translation (MT) metrics which are fine-tuned on human-generated MT quality judgements are robust to domain shifts between training and inference. We find that fine-tuned metrics exhibit a substantial performance drop in the unseen domain scenario relative to metrics that rely on the surface form, as well as pre-trained metrics which are not fine-tuned on MT quality judgments.
@@ 1 Introduction
Given the growing popularity of fine-tuned metrics, it is important to better understand their behavior. Here, we examine the question of domain robustness of fine-tuned metrics. Fine-tuned metrics contain extra parameters on top of the pre-trained model which are initialized randomly (or to zero) and then fine-tuned on human-generated MT quality annotations. The primary source of those annotations is from prior WMT metrics shared tasks, and domains in WMT are often carried over from year to year (e.g. news). This raises the question: are fine-tuned metrics in fact robust across any domain (including domains not seen in training)? Or can their apparent strong performance be attributed in part to the artificially good domain match between training and test data?
@@
Next, we examine how different types of metrics perform on our new bio test set relative to the WMT test set. We find that fine-tuned metrics have substantially lower correlation with human judgments in the bio domain, despite other types of metrics having higher correlation in the bio domain (see Figure 1), indicating they struggle with the training/inference domain mismatch. Finally, we present analysis showing that such performance gap persists throughout different stages of the fine-tuning process and is not the result of a deficiency with the pre-trained model.
@@
Table 1: Metric types considered in this work. The [Uncaptioned image] components have trainable parameters while [Uncaptioned image] use handcrafted heuristics or algorithms and [Uncaptioned image] decodes from a language model. The ref input omitted in the case of reference-free metrics (i.e. quality estimation).
@@ 2 Related Work
Table 1 summarizes the different types of metrics that are commonly used to evaluate MT. The earliest type of MT metrics are Surface-Form metrics, which are purely heuristic and use word- or character-based features. We consider three common Surface-Form metrics: BLEU (Papineni et al., 2002), TER (Snover et al., 2006) and ChrF (Popović, 2015). Metrics like Comet (Rei et al., 2020), Bleurt (Sellam et al., 2020), and UniTE (Wan et al., 2022) start with a pre-trained language model and fine-tune it on human-generated MT quality judgments. We denote these metrics Pre-trained+Fine-tuned. [^2] The WMT metrics task calls these “trained” metrics. Another class of metrics also start with a pre-trained model but do not perform fine-tuning. Examples of such metrics include Prism (Thompson and Post, 2020a, b), which uses the perplexity of a neural paraphraser, and BERTScore (Sun et al., 2022), based on word embedding similarity. We denote such metrics Pre-trained+Algorithem metrics. More recently, metrics like GEMBA (Kocmi and Federmann, 2023) and AutoMQM (Fernandes et al., 2023) have proposed prompting a large language model. We denote these as Pre-trained+Prompt metrics.
@@
Domain specificity for MT metrics was first explored by C. de Souza et al. (2014) for Surface-Form metrics. Sharami et al. (2023) brought attention to the issue of domain adaptation for quality estimation (QE), offering solutions with curriculum learning and generating synthetic scores similar to Heo et al. (2021), Baek et al. (2020), and Zouhar et al. (2023). Sun et al. (2022) examined general-purpose natural language generation metrics and documented their bias with respect to social fairness. For word-level QE, Sharami et al. (2023) reported the lack of robustness of neural metrics.
@@ 3 New Bio MQM Dataset
We create and release new translations and MQM annotations for the system submissions from 21 participants to the WMT21 bio translation shared task (Yeganova et al., 2021). To explore how different the bio domain is from the WMT22 metric task domains, we computed the vocabulary overlap coefficient between each domain. Bio had the smallest average overlap with the WMT domains (0.436) compared to 0.507, 0.486, 0.507, and 0.582 for e-commerce, news, social, and conversation respectively. See Appendix A for full details and example sentences from each domain.
@@ 3.1 Dataset Creation
Step 1: Reference re-translation.
Step 2: Reference quality.
Step 3: MQM annotations.
Finally, we conduct the main MQM annotation on the references and shared task system outputs. In this step, a single annotator rates all translations of a given document (from all systems and the reference). [^3] This allows us to distribute annotation jobs to multiple annotators while still allowing the annotator to access document-level context and ensuring that the whole document is ranked consistently. Our MQM schema follows Freitag et al. (2021) except that we add we add a Critical severity (assigned the same score as Major for backward compatibility). Full annotator instructions are in Appendix D.
@@ 4 Analysis
Figure 2: Segment-level correlation (Kendall’s $\tau$) between metrics and human judgments on the WMT and bio domain. Pre-trained+Fine-tuned metrics have lower correlation on bio than on WMT, while Surface-Form and Pre-trained+Algorithm tend to have higher correlation. Note that there is no public Comet 22 MQM model.
Here we summarize several experiments aimed at understanding the domain gap observed in Pre-trained+Fine-tuned metrics. Full details are provided in Appendix E and our released code.
Fine-tuned metrics exhibit a performance gap on the unseen domain.
We find (Figure 2) that Pre-trained+Fine-tuned metrics exhibit lower correlation with human judgments on the bio MQM test set than the WMT test set, while both Pre-trained+Algorithem and Surface-Form metrics tend to have higher correlations on bio MQM. This gap suggests that fine-tuned metrics struggle with unseen domains, relative to other metric types. We also observe a very large performance gap for Pre-trained+Prompt metrics. Unfortunately, these metrics rely on closed-source LLMs without published training procedures, so we do not know if the underlying LLMs were trained on WMT metric data.
Domain gap persists throughout the fine-tuning process.
@@
Comet is trained in two steps, first on direct assessment (DA) quality annotations and then on MQM annotations, both from WMT shared tasks. In contrast to catastrophic forgetting Goodfellow et al. (2013); Thompson et al. (2019a, b), where a model starts with good general-domain performance and then overfits while being adapted to a new task or domain, we do not see a sharp dropoff in the bio domain when training on more DA/MQM data. This indicates that the model is a poor bio metric at all stages, as opposed to learning and then forgetting.

### v2

@@ Abstract
We introduce a new, extensive multidimensional quality metrics (MQM) annotated dataset covering 11 language pairs in the biomedical domain. We use this dataset to investigate whether machine translation (MT) metrics which are fine-tuned on human-generated MT quality judgements are robust to domain shifts between training and inference. We find that fine-tuned metrics exhibit a substantial performance drop in the unseen domain scenario relative to both metrics that rely on the surface form and pre-trained metrics that are not fine-tuned on MT quality judgments.
@@ 1 Introduction
Given the growing popularity of fine-tuned metrics, it is important to better understand their behavior. Here, we examine the question of domain robustness of fine-tuned metrics. Fine-tuned metrics contain extra parameters on top of the pre-trained model which are initialized randomly (or to zero) and then fine-tuned on human-generated MT quality annotations. The primary source of those annotations is prior WMT metrics shared tasks, and domains in WMT are often carried over from year to year (e.g. news). This raises the question: are fine-tuned metrics in fact robust across any domain (including domains not seen in training)? Or can their apparent strong performance be attributed in part to the artificially good domain match between training and test data?
@@
Next, we examine how different types of metrics perform on our new bio test set relative to the WMT test set. We find that fine-tuned metrics have substantially lower correlation with human judgments in the bio domain, despite other types of metrics having higher correlation in the bio domain (see Figure 1), indicating they struggle with the training/inference domain mismatch. Finally, we present analysis showing that this performance gap persists throughout different stages of the fine-tuning process and is not the result of a deficiency with the pre-trained model.
@@
Table 1: Metric types considered in this work. The [Uncaptioned image] components have trainable parameters while [Uncaptioned image] use handcrafted heuristics or algorithms and [Uncaptioned image] decodes from a language model. The ref input is omitted in the case of reference-free metrics (i.e. quality estimation).
@@ 2 Related Work
Table 1 summarizes the different types of metrics that are commonly used to evaluate MT. The earliest type of MT metrics are Surface-Form metrics, which are purely heuristic and use word- or character-based features. We consider three common Surface-Form metrics: BLEU (Papineni et al., 2002), TER (Snover et al., 2006) and ChrF (Popović, 2015). Metrics like Comet (Rei et al., 2020), Bleurt (Sellam et al., 2020), and UniTE (Wan et al., 2022) start with a pre-trained language model and fine-tune it on human-generated MT quality judgments. We denote these metrics Pre-trained+Fine-tuned. [^2] The WMT metrics task calls these “trained” metrics. Another class of metrics also start with a pre-trained model but do not perform fine-tuning. Examples of such metrics include Prism (Thompson and Post, 2020a, b), which uses the perplexity of a neural paraphraser, and BERTScore (Sun et al., 2022), which is based on cosine similarity of word embeddings. We denote such metrics Pre-trained+Algorithm metrics. More recently, metrics like GEMBA (Kocmi and Federmann, 2023) and AutoMQM (Fernandes et al., 2023) have proposed prompting a large language model. We denote these as Pre-trained+Prompt metrics.
@@
Domain specificity for MT metrics was first explored by C. de Souza et al. (2014) for Surface-Form metrics. Sharami et al. (2023) brought attention to the issue of domain adaptation for quality estimation (QE), offering solutions based on curriculum learning and generating synthetic scores similar to Heo et al. (2021), Baek et al. (2020), and Zouhar et al. (2023). Sun et al. (2022) examined general-purpose natural language generation metrics and documented their bias with respect to social fairness. For word-level QE, Sharami et al. (2023) reported the lack of robustness of neural metrics.
@@ 3 New Bio MQM Dataset
We create and release new translations and MQM annotations for the system submissions from 21 participants to the WMT21 biomedical translation shared task (Yeganova et al., 2021). To explore how different the bio domain is from the WMT22 metric task domains, we computed the vocabulary overlap coefficient between each domain. Bio had the smallest average overlap with the WMT domains (0.436) compared to 0.507, 0.486, 0.507, and 0.582 for e-commerce, news, social, and conversation, respectively. See Appendix A for full details and example sentences from each domain.
Figure 2: Gains in segment-level correlation (Kendall’s $\tau$) when comparing Surface-Form metrics (average performance of BLEU, ChrF, and TER) to a given metric, on the WMT and bio test sets. Gains for Pre-trained+Fine-tuned metrics are much smaller in the unseen bio domain than the WMT domain. Pre-trained+Algorithm metrics, which do not train on prior WMT data, do not exhibit the same bias. See Appendix F for results in tabular form.
@@ 3.1 Dataset Creation
Step 1: Reference re-translation.
Step 2: Reference quality.
Step 3: MQM annotations.
Finally, we conduct the main MQM annotation on the references and shared task system outputs. In this step, a single annotator rates all translations of a given document (from all systems and the reference). [^3] This allows us to distribute annotation jobs to multiple annotators while still allowing the annotator to access document-level context and ensuring that the whole document is ranked consistently. Our MQM schema follows Freitag et al. (2021) except that we add a Critical severity (assigned the same score as Major for backward compatibility). Full annotator instructions are in Appendix D.
@@ 4 Analysis
4.1 Are fine-tuned metrics robust across domains?
Measuring domain robustness.
The performance of a MT metric is typically measured by a certain meta-evaluation metric, such as segment-level Kendall’s $\tau$ correlation with human judgments. Intuitively, one could simply measure domain robustness by comparing the performance of a certain metric on domain A and domain B. This, however, is not straightforward with meta-evaluations for metrics, since performance measured by those meta-evaluations is also affected by factors such as the quantity and quality of the translations included in the dataset, which is often hard to control for.
As a result, we resort to comparisons of relative performance measured against a domain-invariant baseline. To establish such comparison, we make two assumptions:
  1. 1.
    We assume Surface-Form metrics can serve as a domain-invariant baseline, as they are purely based on heuristics and do not involve parameters specifically tuned on a certain domain. We use average performance of BLEU, ChrF, and TER as the baseline to minimize the impact of specific choice of heuristics.
  2. 2.
    We assume segment-level Kendall’s $\tau$ correlation with human judgments has a linear relationship with the objective performance of a metric. Hence, relative performance can be measured by simple linear subtraction.
Observations.
Compared to Surface-Form metrics, we find that Pre-trained+Fine-tuned metrics provide a substantially smaller (sometimes even negative) improvement in human correlation in the bio domain than the WMT domain (see Figure 2). On the other hand, Pre-trained+Algorithm metrics, which have not been trained on WMT data, do not exhibit the same gap. This gap suggests that fine-tuned metrics struggle with unseen domains.
We also observe a very large performance gap for Pre-trained+Prompt metrics. Unfortunately, these metrics rely on closed-source LLMs without published training procedures, so we do not know what data the underlying LLMs were trained on.
4.2 How does fine-tuning affect domain robustness?
Model description.
For this section, we focus on Comet (reference-based) and Comet-QE (reference-free) as they are among the most commonly used MT metrics. The Comet model works by representing the source, the hypothesis and the reference as three fixed-width vectors using a language model, such as XLM-Roberta-large Conneau et al. (2019). These vectors and their combinations serve as an input to a simple feed-forward regressor which is fine-tuned to minimize the MSE loss with human MQM scores. A Comet model is trained in two stages, first on direct assessment (DA) quality annotations and then on MQM annotations, both from WMT shared tasks.
Setup.
We limit our experiments to the En-De, Zh-En and Ru-En language directions because of WMT MQM availability. We largely followed the training recipe in the Comet Github repo555github.com/Unbabel/COMET/tree/master/configs. For details, please refer to our code.
There is high inter-annotator variance in the WMT and bio MQM data. Training on the raw MQM scores is very unstable and therefore per-annotator z-normalizing is necessary to replicate our setup. Note that the publicly available WMT MQM data are not z-normalized.
Observation 1: Domain gap persists throughout the fine-tuning process.
@@ 4.2 How does fine-tuning affect domain robustness?
We would like to understand which stage among the two training stages for Comet accounts for the domain gap. To this end, we retrained Comet with varying epochs on DA/MQM data, shown in Table 4.2. In contrast to catastrophic forgetting Goodfellow et al. (2013); Thompson et al. (2019a, b), where a model starts with good general-domain performance and then overfits while being adapted to a new task or domain, we do not see a sharp dropoff in the bio domain performance when training on more WMT (DA and/or MQM) data. This indicates that the model is a weak bio metric at all stages, as opposed to first learning and then forgetting.
Observation 2: In-domain data dramatically improves Comet.
Figure 3: Average performance (8 seeds) of Comet fine-tuned on varying amounts of MQM bio data.
Generally, including bio MQM annotations in training improves Comet’s performance in the bio test set, increasing correlation from 0.287 to 0.328 with 6k bio judgments. Indeed, just 1k judgements improves correlation to 0.313 (see Figure 3). This rules out the possibility that bio is inherently problematic for Comet’s architecture or fine-tuning strategy.
4.3 How does the pre-trained model affect domain robustness?
Comet and BERTScore are both based on XLM-Roberta-large Conneau et al. (2019), allowing us to explore how the same changes to the pre-trained model affect each metric. To see whether improving the underlying pre-trained model improves Pre-trained+Algorithm metrics built on those pre-trained models, we fine-tune XLM-Roberta with data similar to the WMT and bio domain setup, respectively. Similarly, we also investigate how Prism, another Pre-trained+Algorithm metric, is affected with changes to the pre-trained model. We use Prism with the NLLB multilingual MT models (NLLB Team et al., 2022) as they are larger and more recent than the model released with Prism.
Setup.
Our fine-tuning data covers the four languages of interest, namely English, German, Russian, and Chinese (see Appendix E.2 for a detailed data list). Since NLLB is a translation model, we use only parallel data to fine-tune the model. For the XLM-Roberta case, note that it was fine-tuned with two objectives: masked language model (MLM) and translation language model (TLM). We use both parallel and monolingual data for MLM training and parallel data for TLM training.
[…]
Figure 5: Multiple NLLB MT models are used as the base model for PrismSrc. Fine-tuning the underlying MT model improves the metric. Compute constraints preclude finetuning NLLB-3.3B.
Observations: XLM-Roberta.
For both domains, improving the pre-trained model improves BERTScore but not Comet (see Figure 4). This indicates that the limiting factor for the poor performance of Comet on bio is the effect from its various fine-tuning stages (discussed in Section 4.2), not an underlying weakness in the pre-trained model on bio.
Observations: NLLB.
Our findings are shown in Figure 5. In general, we found that improving the pre-trained models performance (as measured by BLEU on a held out test set) also improved Prism’s performance.
5 Conclusion and Future Work
This paper investigated the performance of machine translation metrics across divergent domains. To this end, we introduced a new, extensive MQM-annotated dataset covering 11 language pairs in the bio domain. Our analysis showed that Pre-trained+Fine-tuned metrics (i.e. those that use prior human quality annotations of MT output) exhibit a larger gap between in-domain and out-of-domain performance than Pre-trained+Algorithm metrics (like BERTScore). Further experiments showed that this gap can be attributed to the DA and MQM fine-tuning stage.
Despite the gap between in-domain and out-of-domain performance, Comet is still the best performing metric on the bio domain in absolute terms. Thus, our findings suggest potential directions for future work including collecting more diverse human judgments for Pre-trained+Fine-tuned metrics and exploring ways to improve the generalization of such metrics during fine-tuning.
Limitations
Our findings are dependent on two empirical assumptions we discussed in section 4.1. To the best of our knowledge, those assumptions are necessary to achieve a fair comparison of metrics across domains, but conclusions may change if our assumptions are refuted in future studies.
We draw conclusions based on a single unseen domain (biomedical). While additional domains would have been preferable, data collection was cost prohibitive.
Context has been shown to be beneficial in machine translation evaluation Läubli et al. (2018); Toral (2020) and some metrics used in this work have document-level versions Vernikos et al. (2022). However, in order to draw fair comparisons with existing metrics which do not yet have a document-level version, we only evaluated metrics at the sentence level.
We focused on segment-level evaluation and did not attempt system-level comparisons because of the limited number of system submissions to the WMT biomedical translation shared task.
Acknowledgements
We would like to thank Georgiana Dinu, Marcello Federico, Prashant Mathur, Stefano Soatto, and other colleagues for their feedback at different stages of drafting.
Ethical Considerations
Our human annotations were conducted through a vendor. Annotators were compensated in accordance to the industry standard – specifically, in the range of \$27.50 to \$37.50 on an hourly basis, depending on the experience of the annotator.
Appendix A Domain Overlap Between WMT and bio
To evaluate the overlap between the WMT and bio domains, we calculate the vocabulary overlap coefficient $(\frac{|A\cap B|}{\min(|A|,|B|)})$ between our new bio MQM dataset and the domains used in the WMT22 metrics shared task. The per-domain overlap matrix is shown in Figure 6. Randomly selected sentences from each domain are provided for illustration in Figure 7.
e-commerce news social conversation biomedical
e-commerce 1.000 0.349 0.511 0.662 0.369
news 0.349 1.000 0.517 0.592 0.359
social 0.511 0.517 1.000 0.494 0.462
conversation 0.662 0.592 0.494 1.000 0.554
biomedical 0.369 0.359 0.462 0.554 1.000
Figure 6: Vocabulary overlap coefficient between the English source-side data for each domain in the WMT22 and our bio dataset.
e-commerce This was one of the first albums I purchased of Keith’s "back in the day".
news Sean Combs has been variously known as Puff Daddy, P. Diddy or Diddy, but this year announced his preference for the names Love and Brother Love.
social The comment about boiling being inefficient is probably correct bc even though the water heater is running continuously, that thing has SO MUCH insulation.
conversation Let me know if you were able to create your new password and sign in with it
biomedical Though neither perfectly sensitive nor perfectly specific for trachoma, these signs have been essential tools for identifying populations that need interventions to eliminate trachoma as a public health problem.
Figure 7: Randomly selected English example sentences from each domain in the WMT22 metrics shared task as well as our new bio dataset.
Appendix B Corpus Statistics
Table 4 shows the size per language pair of our bio MQM dataset, as well as the WMT MQM dataset for comparison. The bio MQM dataset contains roughly 25k annotated segments, covering 11 language pairs. We split the data into test (roughly 75%) and development (roughly 25%) sets.
WMT Bio
Langs Test Train Test Dev Total
De-En - - 2457 903 3360
En-De 18k 28k 2695 917 3612
Es-En - - 1013 309 1322
En-Es - - 1112 330 1442
Ru-En - - 1324 388 1712
En-Ru 19k 16k 825 237 1062
Fr-En - - 1108 352 1460
En-Fr - - 1228 308 1536
Zh-En 23k 27k 2838 913 3751
En-Zh - - 3900 1200 5100
Pt-En - - 701 222 924
All 60k 71k 19k 6k 25k
Table 4: Data split of the bio MQM data released in this work, and WMT22 MQM (Freitag et al., 2022) data. All test results are reported with the test split which is approximately $75\%$ of total. Splits were created to respect document-level boundaries. For WMT, 2022 is used for testing and 2020 and 2021 for training.
Appendix C Translator/Annotator Qualifications
There were 2-4 MQM annotators for each language pair, and a total of 46 annotators. All linguists had experience in translating/post-editing/reviewing content in the bio domain. This was the main requirement to be able to work on the project. The other qualification criteria for this project were in line with the ISO standard 17100. In particular, the linguists met one or more of the following criteria: (1) A recognized higher education degree in translation; (2) Equivalent third-level degree in another subject plus a minimum of two years of documented professional translation experience; (3) A minimum of five years of documented professional translation experience; (4) Native speaker of the target language. Although linguists were experts in the bio domain, not all of them were experts in MQM annotation. For this reason, the annotators completed an MQM quiz before onboarding them to ensure they understood the guidelines and requirements.
For the translation and post-editing tasks, we used a two step process (initial post editor + reviewer). In each case the reviewer was a linguist with experience translating medical texts. There were no specific educational or vocational stipulations on that medical qualification, however they were asked to provide a medical-text-specific translation test for us to be onboarded for the project. The initial post-editor in each case was a linguistic expert, but not specifically an expert in medical translations, which is why we followed up with reviewers to ensure contents were translated accurately. Linguists had to demonstrate the following to onboard to the project: (1) At least 3+ years of professional translation experience (2) Proven proficiency in English writing skills (3) In-depth understanding and exposure to the language (4) Strong ability in translating, reviewing, adjusting, and providing adaptation for various writing styles of particular requests.
Appendix D MQM Annotation Guidelines
Below, we reproduce the MQM annotation guidelines that we provided to the annotators.
Overview:
You are asked to evaluate the translations using the guidelines below, and assign error categories and severities considering the context segments available.
Task:
  1. 1.
    Please identify all errors within each translated segment, up to a maximum of five.
    1. (a)
      If there are more than five errors, identify only the five most severe.
    2. (b)
      If it is not possible to reliably identify distinct errors because the translation is too badly garbled or is unrelated to the source, then mark a single Unintelligible error that spans the entire segment
    3. (c)
      Annotate segments in natural order, as if you were reading the document. You may return to revise previous segments.
  2. 2.
    To identify an error, highlight the relevant span of text.
    1. (a)
      Omission and Source error should be tagged in the source text.
      1. i.
        All other errors should be tagged in the target text.
    2. (b)
      Unintelligible error should have an entire sentence tagged; if you think a smaller span is needed, then you should select another error category (Mistranslation, etc.).
  3. 3.
    Select a category/sub-category and severity level from the available options.
  4. 4.
    When identifying errors, please be as fine-grained as possible.
    1. (a)
      If a sentence contains more than one error of the same category, each one should be logged separately. For example, if a sentence contains two words that are each mistranslated, two separate mistranslation errors should be recorded.
    2. (b)
      If a single stretch of text contains multiple errors, you only need to indicate the one that is most severe.
      1. i.
        If all have the same severity, choose the first matching category listed in the error typology (e.g. Accuracy, then Fluency, then Terminology, etc.).
    3. (c)
      For repetitive errors that appear systematically through the document: please annotate each instance with the appropriate weight.
  5. 5.
    Please pay particular attention to the context when annotating. You will be shown several context segments before and after the segment for evaluation. If a translation is questionable on its own but is fine in the context of the document, it should not be considered erroneous; conversely, if a translation might be acceptable in some context, but not within the current document, it should be marked as wrong.
Delivery format:
  • file format: a TSV with additional columns for error categories and severity + JSON
    • for multiple errors in one segment: additional row for each error + severity
    • text spans will be highlighted for the annotation process and exported as tag
Error categories:
Table 5
Tag
Location
Accuracy – errors occurring when the target text does not accurately correspond to the propositional content of the source text, introduced by distorting, omitting, or adding to the message Mistranslation Target content that does not accurately represent the source content. Target
Addition Target content that includes content not present in the source. Target
Omission Errors where content is missing from the translation that is present in the source. Source
Untranslated Errors occurring when a text segment that was intended for translation is left untranslated in the target content. Target
Linguistic Conventions (former Fluency) - errors related to the linguistic well-formedness of the text, including problems with, for instance, grammaticality and mechanical correctness. Grammar Error that occurs when a text string (sentence, phrase, other) in the translation violates the grammatical rules of the target language. Target
Punctuation Punctuation incorrect for the locale or style. Target
Spelling Error occurring when the letters in a word in an alphabetic language are not arranged in the normally specified order. Target
Character encoding Error occurring when characters garbled due to incorrect application of an encoding. Target
Register Errors occurring when a text uses a level of formality higher or lower than required by the specifications or by common language conventions. Target
Terminology - errors arising when a term does not conform to normative domain or organizational terminology standards or when a term in the target text is not the correct, normative equivalent of the corresponding term in the source text.
Inconsistent use of terminology Use of multiple terms for the same concept (technical terms, medical terms, etc.) Target
Wrong term Use of term that it is not the term a domain expert would use or because it gives rise to a conceptual mismatch. Target
Style Non-fluent Text does not sound fluent or natural as if it were translated by a non-native speaker or because the translation is following the source too closely. Target
Locale Conventions - errors occurring when the translation product violates locale-specific content or formatting requirements for data elements. Number format Target
Currency format Target
Measurement format Target
Time format Target
Date format Target
Address format Target
Telephone format Target
Other any error that does not fit the categories above Target
Source errors source error The error that occurs in the source. All source errors (e.g. non-fluent source) should be annotated as source errors — no sub-categories need to be selected. If the source error caused a target error: - if the source error and target errors belong to the same category, then only flag the source. -If source and target errors belong to different categories - even if you know that the source error caused the translation error - do flag both. Source
Unintelligible So many errors, or errors are so outrageous, that text becomes incomprehensible, and it is hard to pinpoint a specific error type. Target. Tag the entire sentence. If the span is smaller, then a different category should be applied, such as Mistranslation, Untranslated, etc.
Table 5: MQM error categories provided in annotator instructions.
Severity (no weights, just severity):
Table 6
severity Definition Source example Translation example
Neutral Neutral issues are items that need to be noted for further attention or fixing but which should not count against the translation. This severity level can be perceived as a flag for attention that does not impose a penalty. It should be used for “preferential errors” (i.e, items that are not wrong, per se, but where the reviewer or requester would like to see a different solution). Source: Join us in celebrating 10 years of the company! Target: Join us to celebrate 10 years of the company!
Minor Minor issues are issues that do not impact usability or understandability of the content. If the typical reader/user is able to correct the error reliably and it does not impact the usability of the content, it should be classified as minor. S1: Accurately distinguish between legitimate and high-risk account registrations S2: See how organizations worldwide are using fraud detection. T1: Accurately distinguish between legitimate and high- risk account registrations T2: See how organization worldwide are using fraud detection.
Major errors that would impact usability or understandability of the content but which would not render it unusable. For example, a misspelled word that may require extra effort for the reader to understand the intended meaning but does not make it impossible to comprehend should be labeled as a major error. Additionally, if an error cannot be reliably corrected by the reader/user (e.g., the intended meaning is not clear) but it does not render the content unfit for purpose, it should be categorized as major. Source: Set the performance to 50 percent Target: Set performance 50 percent
Critical errors that would render a text unusable, which is determined by considering the intended audience and specified purpose. For example, a particularly bad grammar error that changes the meaning of the text would be considered Critical. Critical errors could result in damage to people, equipment, or an organization’s reputation if not corrected before use. If the error causes the text to become unintelligible, it would be considered Critical. S1: Set the device on the highest temperature setting. S2: The next step would be to identify the point of leakage. S3: 1.3 degrees T1: Set the device on the lowest temperature setting. T2: It would be to identify the next point of leakage. T3: 1,300 degrees
Table 6: Severity examples and explanations provided in MQM annotation instructions.
Appendix E Supplementary Information on Experiments
E.1 Training Steps and Compute Time for Experiments
The overall training consists of the following steps (compute times using a single A10 GPU). The times are per epoch and some experiments require training for multiple epochs.
  • Language modeling $\rightarrow$ XLM-Roberta, 10hr/ep.
  • DA scores regression $\rightarrow$ CometDA, 10hr/ep.
  • MQM scores regression $\rightarrow$ Comet, 1hr/ep.
E.2 List of Data for Fine-Tuning Pre-Trained Model
For WMT domain, we used news-commentary v18.1 dataset666data.statmt.org/news-commentary/v18.1/ for all languages. For the bio domain, we list the data in Table 7.
Data Type Language(s) Dataset Lines
Parallel en-de
UFAL Medical Corpus
Yeganova et al. (2021)
3M
en-de MEDLINE Yeganova et al. (2021) 35k
en-ru 29k
en-zh 19k
Monoling. En CORD Wang et al. (2020) 1M
De
Animal Experiments777www.openagrar.de/receive/openagrar_mods_00046540?lang=en
GERNERMED
Frei and Kramer (2023)
250k
Ru Medical QA 250k
Zh Chinese Medical Dataset888huggingface.co/datasets/shibing624/medical 2M
Table 7: Collection of bio domain data used in pre-trained model fine-tuning experiments.
Appendix F Raw Scores for Figure 2
The segment-level correlation (Kendall’s $\tau$) scores used to compute improvements in Figure 2 are provided in Table 8. Note that there is no public Comet 22 MQM model.
Type Metric Test:WMT Test:Bio
Surface-Form BLEU 0.134 0.213
ChrF 0.151 0.192
TER 0.140 0.100
Pre-trained+Algorithm PRISMREF 0.216 0.242
PRISMSRC 0.121 0.267
BERTScore 0.216 0.227
Pre-trained+Prompt GEMBADAV3 0.280 0.159
GEMBADAV3.QE 0.222 0.173
Pre-trained+Fine-tuned COMETMQM.21 0.328 0.249
COMETQE.21 0.294 0.205
COMETDA.21 0.309 0.284
COMETINHO.21 0.255 0.182
COMETDA.22 0.304 0.269
UniTE 0.301 0.249
BLEURT 0.214 0.100
Table 8: Segment-level correlation (Kendall’s $\tau$) between metrics and human judgments on the WMT and bio domain. Pre-trained+Fine-tuned metrics have lower correlation on bio than on WMT, while Surface-Form and Pre-trained+Algorithm tend to have higher correlation.